300 pennies are tossed into the air. Determine the probability of:
1) P(125 is less than or = X less than or = 170 tails)
2) If your co worker is willing to wager 10.00 on this, should I take the wager and why?

Answer by stanbon(57967)

300 pennies are tossed into the air. Determine the probability of:
1) P(125 is less than or = X less than or = 170 tails)
z(125)=(125-300*(1/2))/sqrt(300(1/2)(1/2))= -2.88675
z(170)=(170-150)/sqrt(300*(1/4))=2.3094
---------
P(125<=x<=170)=P(-2.88675<=z<=2.3094)=0.98759
=================
2) If your co worker is willing to wager 10.00 on this, should I take the wager and why?
No, because the probability it will happen is almost 99%
=========
Cheers,
Stan H.